# Magnetorheological Characterization of Blood Analogues Seeded with Paramagnetic Particles

### v1

@@ 3.2.2 Magnetorheological benchmark measurements
Figure 5 shows the obtained Pipkin diagrams and viscoelastic moduli plots of the benchmark measurements (15 wtof MyOne particles under the effects of a 250 mT magnetic field density). Across the whole frequency/amplitude range the rheological differences between the continuous phases do not appear, as the results of both the seeded NBa and VBa samples are very similar, aligning with the steady shear findings. For the largest tested frequencies, $\omega=1$ and 2 rad/s, there are oscillations in the Lissajous curves that may be due to inertial issues. Compared to the data gathered with the unseeded analogues, there is a significant increase of the viscoelastic moduli across the tested range which allows to escape from the experimental limits, meaning the setup can capture the magnetorheological properties of the seeded samples, at least for this particle type, concentration and applied magnetic field density.
With frequency increase, the elastic response is slightly reduced. Still, within the tested range, the magnetorheological behaviour seems to be practically frequency-independent, which is expected at low frequency17. At low amplitude, $\gamma_{0}=0.1$, the response is elastic-dominated, with elliptical Lissajous and a slight bending of the viscous stress which points toward a degree of non-linearity. With strain increase the response becomes progressively more viscous (moduli crossover between $1\lesssim\gamma_{0}\lesssim 10$) and non-linear. At high strain, $\gamma_{0}=100$, the elastic contribution is practically negligible and the Lissajous curves display a viscoplastic response.
@@
Figure 5: Oscillatory shear data gathered with the NBa (left) and VBa (right) seeded with 15 wtof MyOne particles under a 250 mT magnetic field, with the PP20 MRD P2. (Top) Pipkin diagrams depicting the Lissajous curves. The blue and red curves represent, respectively, the elastic (stress vs strain) and viscous (stress vs strain rate) Lissajous curves. The elastic and viscous stresses are also lightly plotted in the respective colours and the maximum measured stress is presented above each plot. (Bottom) First harmonic loss ($G^{{}^{\prime\prime}}_{1}$, in circular markers) and storage ($G^{{}^{\prime}}_{1}$, in triangular markers) moduli and experimental limits associated with low-torque issues and instrument and sample inertia (corrected for the geometry’s gap-error).
The moduli plots of Figure 5 show, from the decrease of either $G^{\prime}_{1}$ and $G^{\prime\prime}_{1}$, an average elastic softening and viscous thinning, but the intricacies of the response’s non-linearity remain obscure (the Lissajous curves do shed some light on the non-linear response, albeit qualitatively). We used the MITlaos software38 to quantitatively describe the sample’s non-linear magnetorheological response through FT-Chebyshev analysis. Because there seem to be no behavioural differences between the samples with either analogue, nor frequency-dependence apart from seemingly-inertial errors at large frequency, we will focus on the results obtained with the NBa benchmark sample (15 wtof MyOne @ 250 mT) at the minimum driving frequency ($\omega=0.1$ rad/s). The raw stress and strain data was imported to MITlaos and the signal was filtered for the first 15 harmonics. Figure 6 exhibits the obtained viscoelastic moduli relevant in the non-linear regime ($G_{1}^{\prime}$, $G_{M}^{\prime}$, $G_{L}^{\prime}$ and $\eta_{1}^{\prime}$, $\eta_{M}^{\prime}$, $\eta_{L}^{\prime}$), the Chebyshev coefficient spectrum (until the 15th harmonic) with the corresponding reconstructed Lissajous curves, and the relative contribution of the third-order coefficients.
@@
Figure 6: Oscillatory shear data gathered with the NBa with 15 wtof MyOne particles under a 250 mT magnetic field, with the PP20 MRD P2 ($\omega=0.1$ rad/s). Scaled Chebyshev coefficient spectrum, $e_{\mathrm{n}}/e_{1}$ and $v_{\mathrm{n}}/v_{1}$ (with reconstructed Lissajous curves), variation of the scaled third harmonic coefficients, $e_{3}/e_{1}$ and $v_{3}/v_{1}$, and relevant viscoelasticity measures (elastic: $G_{1}^{\prime}$, $G_{M}^{\prime}$, $G_{L}^{\prime}$, and viscous: $\eta_{1}^{\prime}$, $\eta_{M}^{\prime}$, $\eta_{L}^{\prime}$) with applied deformation/deformation rate (negative $G_{M}^{\prime}$ values are presented as open markers). Analysed with MITlaos38.
@@
Considering the elastic measures (left column of Figure 6), at the minimum strain, $\gamma_{0}=0.1$, all elastic moduli approach the same value, $G_{M}^{\prime}\approx G_{L}^{\prime}\approx G_{1}^{\prime}$, relating to a linear elastic response also showcased by the lack of high-order contributions on the elastic Chebyshev spectrum. Increasing the strain leads to a divergence of $G_{M}^{\prime}$ and $G_{L}^{\prime}$, with the first decreasing faster. This onset of elastic non-linearity can also be observed by the appearance of high-order Chebyshev coefficients which become progressively more significant with input-strain increase. Notably, at the maximum imposed deformation ($\gamma_{0}=100$) the minimum strain modulus takes on a negative value, $G_{M}^{\prime}\approx-2.22\times 10^{-3}$ Pa (negative values of $G_{M}^{\prime}$ are presented as open markers), corresponding to a negative slope of the elastic Lissajous at $\gamma=0$. Negative values of $G_{M}^{\prime}$ are associated with secondary loops of the viscous Lissajous curve in a strong non-linear response where the sample is discharging elastic stress at a faster rate than the deformation is being applied, and are expected at large strain and frequency39. Looking at the Lissajous curves in question, there are no prominent secondary loops on the viscous response and the negative slope of the elastic stress appears to be a result of local oscillations of the stress at at $\dot{\gamma}\to\dot{\gamma}_{0}$, which should be due to either inertial or low-torque issues as the sample softens. The third-order Chebyshev coefficient contribution is always positive in the NLVE ($e_{3}>0$), meaning the average elastic softening is being driven by the strain rate.
@@
Regarding the viscous properties (right column of Figure 6), already at the minimum tested rate ($\dot{\gamma}_{0}=0.01$ s-1) $\eta_{L}^{\prime}>\eta_{M}^{\prime}$ and a non-negligible positive contribution of $v_{3}$ is observable, meaning the viscous non-linearity onsets before the elastic. At $\dot{\gamma}_{0}=0.1$ s-1, $\eta_{M}^{\prime}\approx\eta_{L}^{\prime}$ which could point towards a return to linear behaviour, if we had not also evaluated the Lissajous and the Chebyshev coefficients, which clearly display a non-linear response. There is a significant transition between $0.10.01<\dot{\gamma}_{0}<$ 0.1 s-1 (or $10.1<\gamma_{0}<$ 1), with $v_{3}$ changing from positive to negative which points towards significant microstructural changes.
@@
Figure 7: Oscillatory shear data gathered with the NBa with 15 wtof MyOne particles under a 250 mT magnetic field, with the PP20 MRD P2 ($\omega=0.1$ rad/s). (Top) first harmonic loss ($G^{{}^{\prime\prime}}_{1}$, in circular markers) and storage ($G^{{}^{\prime}}_{1}$, in triangular markers) moduli and elastic and viscous Lissajous curves (reconstructed for the first 15 harmonics), in blue and red respectively, and corresponding stresses (lightly plotted). (Bottom) Scaled 3rd-order Chebyshev coefficients, $e_{3}/e_{1}$ and $v_{3}/v_{1}$, and relevant viscoelasticity measures (elastic: $G_{1}^{\prime}$, $G_{M}^{\prime}$, $G_{L}^{\prime}$, and viscous: $\eta_{1}^{\prime}$, $\eta_{M}^{\prime}$, $\eta_{L}^{\prime}$). Data from MITlaos38.
@@
Entering the NLVE, and considering the elastic response first, before any significant contribution from high-order harmonics, a slight decrease of $G^{\prime}_{1}$ is associated to first-harmonic non-linearities22. From $\gamma_{0}\gtrsim 0.1$, the decrease of $G^{\prime}_{1}$ is accelerated, accompanied by an increasing positive contribution of $e_{3}$ and a faster decrease of $G^{\prime}_{M}$. Throughout the entire tested strain range, $G_{1}^{\prime}$ decreases, with $e_{3}>0$ and $G^{\prime}_{M}>G^{\prime}_{L}$, which relates to an elastic softening of the sample driven by the strain rate (lesser elastic stress than the linear basis at $\dot{\gamma}\to\dot{\gamma}_{0}$).
@@
Regarding the viscous non-linearities, In the beginning of the NLVE we observe the viscous non-linearity onseting first, at $\dot{\gamma}_{0}\approx 0.005$ s-1 ($\gamma_{0}\approx 0.05$), with an increase of $v_{3}/v_{1}$ and $\eta^{\prime}_{L}$ that should be associated with the overshoot of $G^{\prime\prime}_{1}$ (type III behaviour18). As such, the viscous non-linearity begins with viscous thickening driven by the strain rate, returning larger viscous stresses than the linear baseline at $\dot{\gamma}\to\dot{\gamma}_{0}$, where the microstructural rearrangements occur. Increasing the strain rate we observe the simultaneous decrease of $G^{\prime\prime}_{1}$, $\eta^{\prime}_{L}$ and $v_{3}/v_{1}$, with the latter switching signs at $\dot{\gamma}_{0}\approx 0.05$ s-1 ($\gamma_{0}\approx 0.5$). This is indicative of a transition from viscous thickening to thinning, both driven by the strain rate, which ought to be associated with the onset of large intracycle structural modifications and, at some point, the beginning of microstructural breakdown at $\dot{\gamma}\to\dot{\gamma}_{0}$.
@@
The peak of $G^{\prime\prime}_{1}$ and $v_{3}/v_{1}$ are simultaneous, meaning the onset of viscous thinning (decrease of $G^{\prime\prime}_{1}$) occurs when the contribution of $v_{3}$ is still positive. This could be due to the non-constant shear rate applied in PP geometries that leads to different microstructural responses along the geometry radius, thus having no significant microstructural relevance, or it could signify a momentary strain-dependence of the non-linearity. Considering the second hypothesis, due to the time-dependent nature of oscillatory shear, it could be possible that for a small strain, chain rotation is sufficient to comply with the applied deformation, independently of the strain rate, given that the oscillatory frequency is small. As such, only beyond a critical strain would the microstructure begin to breakdown due to viscous shearing forces. To further understand the mechanism behind the beginning of viscous thinning, we performed a short numerical work.
@@ 3.3 Numerical analysis
$\boldsymbol{F_{\mathrm{d}}}=3\,\pi\,\eta\,d_{\mathrm{p}}(\boldsymbol{u_{\mathrm{f}}}-\boldsymbol{u_{\mathrm{p}}})\,,$ (6)
$\boldsymbol{F_{\mathrm{a,ij}}}=\frac{3\,\mu_{0}\,\mu_{\mathrm{f}}\,m^{2}}{4\,\pi\,r_{\mathrm{ij}}^{4}}\left[\left(1-5\left(\hat{m}\cdot\hat{r}_{\mathrm{ij}}\right)^{2}\right)\hat{r}_{\mathrm{ij}}+2\left(\hat{m}\cdot\hat{r}_{\mathrm{ij}}\right)\hat{m}\right]\,,$ (7)
$\boldsymbol{F_{\mathrm{r,ij}}}=2\,\frac{3\,\mu_{0}\,\mu_{\mathrm{f}}\,m^{2}}{4\,\pi\,d_{\mathrm{p}}^{4}}\exp\left[{-30\left(\frac{r_{\mathrm{ij}}}{d_{\mathrm{p}}}-1\right)}\right]\hat{r}_{\mathrm{ij}}\,,$ (8)
where $(\boldsymbol{u_{\mathrm{f}}}-\boldsymbol{u_{\mathrm{p}}})$ is the particle velocity ($\boldsymbol{u_{\mathrm{p}}}$) relative to the fluid ($\boldsymbol{u_{\mathrm{f}}}$), $\mu_{0}$ and $\mu_{\mathrm{f}}$ are the vacuum and fluid permeabilities, respectively ($\mu_{\mathrm{f}}=1$), $\boldsymbol{r_{\mathrm{ij}}}$ is the interparticle vector (magnitude $r_{\mathrm{ij}}$ and direction $\hat{r}_{\mathrm{ij}}$) and $\boldsymbol{m_{\mathrm{ij}}}$ is the magnetic moment (with magnitude $m=4/3\pi(d_{\mathrm{p}}/2)^{3}\,\rho_{\mathrm{p}}\,M^{*}$ and direction $\hat{m}_{\mathrm{ij}}$, perpendicular to the flow direction). In Figure 8 is shown a schematic of the forces applied on an element of a 2-particle chain.
@@
Before heading to the oscillatory shear simulations, we tested for steady shear first to evaluate the critical shear rates for chain subdivision and complete breakdown. The shear rate was varied by imposing a top wall velocity $U=\dot{\gamma}\,h_{\mathrm{c}}$. It was found that both chains breakdown to individual particles when $\dot{\gamma}_{\mathrm{brk.}}\approx 0.79$ s-1, which was expected because chain breakdown is equivalent to the subdivision of any two-element chain. On the other hand, for $l_{\mathrm{chain}}=4$ the critical subdivision shear rate was $\dot{\gamma}_{\mathrm{sub.}}\approx 0.23$ s-1, while the largest chain with $l_{\mathrm{chain}}=10$ subdivided earlier, at $\dot{\gamma}_{\mathrm{sub.}}\approx 0.04$ s-1 (data not shown here for conciseness). Still considering steady shear, we can also gather information on the critical angle for chain break, $\phi_{\mathrm{c}}$. It is relevant, however, that while in reality the particle contact force is a step function triggered for $r_{\mathrm{ij}}=d_{\mathrm{p}}$, the excluded-volume force in our formulation is significative for small interparticle distances and a global magnetic force can be computed as $\boldsymbol{F_{\mathrm{m,ij}}}=\boldsymbol{F_{\mathrm{a,ij}}}+\boldsymbol{F_{\mathrm{r,ij}}}$. In Figure 9 are plotted the forces acting on an a 2-element chain subject to a steady flow with shear rate $\dot{\gamma}=1.1$ and 0.7 s-1 against the chain angle, $\phi$. On the left column of Figure 9 are shown the global magnetic and viscous drag forces and on the right column the magnetic attraction and repulsive excluded-volume forces are shown decoupled (see Figure 8 for reference).
@@
We can see that the viscous shearing force weakens slightly as the chain rotates due to the reduction of the vertical disparity between the particles ($\Delta y=r_{\mathrm{ij}}\,\cos(\phi)$). Contrarily, the magnetic force increases on both axis ($x$ and $y$) as the magnetic dipole interactions try to re-align the chain with the magnetic field (aligned with $y$). If this simultaneous decrease of the relative drag and increase of the magnetic force is sufficient to reach an equilibrium ($F_{\mathrm{d,x}}+F_{\mathrm{m,x}}=0$) the chain remains stable, as is the case for $\dot{\gamma}=0.7$ s-1. However, if, despite the force convergence, the viscous drag remains too strong (the case where $\dot{\gamma}=1.1$ s-1), the chain breaks. Computing the minimum of the force difference $F_{\mathrm{d,x}}+F_{\mathrm{m,x}}$, we reach the critical chain angle $\phi_{\mathrm{c}}\approx 33.5^{\circ}$ (plotted as a vertical dotted line). With the current method, however, the critical angle is slightly dependent on the applied shear rate and the employed time step. We understand that this is due to the excluded-volume force feeding instantaneous distancing arisen from overpowering viscous forces, which inhibits pure chain rotation and, therefore, contributes to a premature chain break. As can be seen from the plots of the decoupled magnetic attraction, the excluded-volume force reduces the chain’s stability. Still, it also delays the critical angle, as the minimum of the function $F_{\mathrm{d,x}}+F_{\mathrm{a,x}}$ is found slightly earlier, at $\phi_{\mathrm{c}}\approx 28.3^{\circ}$.

### v2

@@ 3.2.2 Magnetorheological benchmark measurements
Figure 5 shows the obtained Pipkin diagrams and viscoelastic moduli plots of the benchmark measurements (15 wt% of MyOne particles under the effects of a 250 mT magnetic field density). Across the whole frequency/amplitude range the rheological differences between the continuous phases do not appear, as the results of both the seeded NBa and VBa samples are very similar, aligning with the steady shear findings. For the largest tested frequencies, $\omega=1$ and 2 rad/s, there are oscillations in the Lissajous curves that may be due to inertial issues. Compared to the data gathered with the unseeded analogues, there is a significant increase of the viscoelastic moduli across the tested range which allows to escape from the experimental limits, meaning the setup can capture the magnetorheological properties of the seeded samples, at least for this particle type, concentration and applied magnetic field density.
With frequency increase, the elastic response is slightly reduced. Still, within the tested range, the magnetorheological behaviour seems to be practically frequency-independent, which is expected at low frequencyparthasarathy1999. At low amplitude, $\gamma_{0}=0.1$, the response is elastic-dominated, with elliptical Lissajous and a slight bending of the viscous stress which points toward a degree of non-linearity. With strain increase the response becomes progressively more viscous (moduli crossover between $1\lesssim\gamma_{0}\lesssim 10$) and non-linear. At high strain, $\gamma_{0}=100$, the elastic contribution is practically negligible and the Lissajous curves display a viscoplastic response.
@@
Figure 5: Oscillatory shear data gathered with the NBa (left) and VBa (right) seeded with 15 wt% of MyOne particles under a 250 mT magnetic field, with the PP20 MRD P2. (Top) Pipkin diagrams depicting the Lissajous curves. The blue and red curves represent, respectively, the elastic (stress vs strain) and viscous (stress vs strain rate) Lissajous curves. The elastic and viscous stresses are also lightly plotted in the respective colours and the maximum measured stress is presented above each plot. (Bottom) First-harmonic loss ($G^{\prime\prime}_{1}$, in circular markers) and storage ($G^{\prime}_{1}$, in triangular markers) moduli and experimental limits associated with low-torque issues and instrument and sample inertia (corrected for the geometry’s gap-error).
The moduli plots of Figure 5 show, from the decrease of either $G^{\prime}_{1}$ and $G^{\prime\prime}_{1}$, an average elastic softening and viscous thinning, but the intricacies of the response’s non-linearity remain obscure (the Lissajous curves do shed some light on the non-linear response, albeit qualitatively). We used the MITlaos softwaremitlaos to quantitatively describe the sample’s non-linear magnetorheological response through FT-Chebyshev analysis. Because there seem to be no behavioural differences between the samples with either analogue, nor frequency-dependence apart from seemingly-inertial errors at large frequency, we will focus on the results obtained with the NBa benchmark sample (15 wt% of MyOne under 250 mT) at the minimum driving frequency ($\omega=0.1$ rad/s). The raw stress and strain data were imported to MITlaos, and the signal was filtered for the first 15 harmonics. Figure 6 exhibits the obtained viscoelastic moduli relevant in the non-linear regime ($G_{1}^{\prime}$, $G_{M}^{\prime}$, $G_{L}^{\prime}$ and $\eta_{1}^{\prime}$, $\eta_{M}^{\prime}$, $\eta_{L}^{\prime}$), the Chebyshev coefficient spectrum (until the 15th harmonic) with the corresponding reconstructed Lissajous curves, and the relative contribution of the third-order coefficients.
@@
Figure 6: Oscillatory shear data gathered with the NBa with 15 wt% of MyOne particles under a 250 mT magnetic field, with the PP20 MRD P2 ($\omega=0.1$ rad/s). Scaled Chebyshev coefficient spectrum, $e_{\mathrm{n}}/e_{1}$ and $v_{\mathrm{n}}/v_{1}$ (with reconstructed Lissajous curves), variation of the scaled third harmonic coefficients, $e_{3}/e_{1}$ and $v_{3}/v_{1}$, and relevant viscoelasticity measures (elastic: $G_{1}^{\prime}$, $G_{M}^{\prime}$, $G_{L}^{\prime}$, and viscous: $\eta_{1}^{\prime}$, $\eta_{M}^{\prime}$, $\eta_{L}^{\prime}$) with applied deformation/deformation rate (negative $G_{M}^{\prime}$ values are presented as open markers). Analysed with MITlaosmitlaos.
@@
Considering the elastic measures (left column of Figure 6), at the minimum strain, $\gamma_{0}=0.1$, all elastic moduli approach the same value, $G_{M}^{\prime}\approx G_{L}^{\prime}\approx G_{1}^{\prime}$, relating to a linear elastic response also showcased by the lack of high-order contributions on the elastic Chebyshev spectrum. Increasing the strain leads to a divergence of $G_{M}^{\prime}$ and $G_{L}^{\prime}$, with the first decreasing faster. This onset of elastic non-linearity can also be observed by the appearance of high-order Chebyshev coefficients, which become progressively more significant with input-strain increase. Notably, at the maximum imposed deformation ($\gamma_{0}=100$), the minimum strain modulus takes on a negative value, $G_{M}^{\prime}\approx-2.22\times 10^{-3}$ Pa (negative values of $G_{M}^{\prime}$ are presented as open markers), corresponding to a negative slope of the elastic Lissajous at $\gamma=0$. Negative values of $G_{M}^{\prime}$ are associated with secondary loops of the viscous Lissajous curve in a strong non-linear response where the sample is discharging elastic stress at a faster rate than the deformation is being applied, and are expected at large strain and frequencyewoldt2010. Looking at the Lissajous curves in question, there are no prominent secondary loops on the viscous response and the negative slope of the elastic stress appears to be a result of local oscillations of the stress at $\dot{\gamma}\to\dot{\gamma}_{0}$, which should be due to either inertial or low-torque issues as the sample softens. The third-order Chebyshev coefficient contribution is always positive in the NLVE ($e_{3}>0$), meaning the strain rate drives the average elastic softening.
@@
Regarding the viscous properties (right column of Figure 6), already at the minimum tested rate ($\dot{\gamma}_{0}=0.01$ s-1) $\eta_{L}^{\prime}>\eta_{M}^{\prime}$ and a non-negligible positive contribution of $v_{3}$ is observable, meaning the viscous non-linearity onsets before the elastic. At $\dot{\gamma}_{0}=0.1$ s-1, $\eta_{M}^{\prime}\approx\eta_{L}^{\prime}$, which could point towards a return to linear behaviour had we not also evaluated the Lissajous and the Chebyshev coefficients, which clearly display a non-linear response. There is a significant transition between $0.01<\dot{\gamma}_{0}<0.1$ s-1 (or $0.1<\gamma_{0}<1$), with $v_{3}$ changing from positive to negative, which points towards significant microstructural changes.
@@
Figure 7: Oscillatory shear data gathered with the NBa with 15 wt% of MyOne particles under a 250 mT magnetic field, with the PP20 MRD P2 ($\omega=0.1$ rad/s). (Top) first-harmonic loss ($G^{\prime\prime}_{1}$, in circular markers) and storage ($G^{\prime}_{1}$, in triangular markers) moduli and elastic and viscous Lissajous curves (reconstructed for the first 15 harmonics), in blue and red respectively, and corresponding stresses (lightly plotted). (Bottom) Scaled 3rd-order Chebyshev coefficients, $e_{3}/e_{1}$ and $v_{3}/v_{1}$, and relevant viscoelasticity measures (elastic: $G_{1}^{\prime}$, $G_{M}^{\prime}$, $G_{L}^{\prime}$, and viscous: $\eta_{1}^{\prime}$, $\eta_{M}^{\prime}$, $\eta_{L}^{\prime}$). Data from MITlaosmitlaos.
@@
Entering the NLVE, and considering the elastic response first, before any significant contribution from high-order harmonics, a slight decrease of $G^{\prime}_{1}$ is associated to first-harmonic non-linearitiesewoldt2013. From $\gamma_{0}\gtrsim 0.1$, the decrease of $G^{\prime}_{1}$ is accelerated, accompanied by an increasing positive contribution of $e_{3}$ and a faster decrease of $G^{\prime}_{M}$. Throughout the entire tested strain range, $G_{1}^{\prime}$ decreases, with $e_{3}>0$ and $G^{\prime}_{M}>G^{\prime}_{L}$, which relates to an elastic softening of the sample driven by the strain rate (lesser elastic stress than the linear basis at $\dot{\gamma}\to\dot{\gamma}_{0}$).
@@
Regarding the viscous non-linearities, In the beginning of the NLVE we observe the viscous non-linearity onseting first, at $\dot{\gamma}_{0}\approx 0.005$ s-1 ($\gamma_{0}\approx 0.05$), with an increase of $v_{3}/v_{1}$ and $\eta^{\prime}_{L}$ that should be associated with the overshoot of $G^{\prime\prime}_{1}$ (type III behaviourhyun2002). As such, the viscous non-linearity begins with viscous thickening driven by the strain rate, returning larger viscous stresses than the linear baseline at $\dot{\gamma}\to\dot{\gamma}_{0}$, where the microstructural rearrangements occur. Increasing the strain rate we observe the simultaneous decrease of $G^{\prime\prime}_{1}$, $\eta^{\prime}_{L}$ and $v_{3}/v_{1}$, with the latter switching signs at $\dot{\gamma}_{0}\approx 0.05$ s-1 ($\gamma_{0}\approx 0.5$). This is indicative of a transition from viscous thickening to thinning, both driven by the strain rate, which ought to be associated with the onset of large intracycle structural modifications and, at some point, the beginning of microstructural breakdown at $\dot{\gamma}\to\dot{\gamma}_{0}$.
@@
The peak of $G^{\prime\prime}_{1}$ and $v_{3}/v_{1}$ are simultaneous, meaning the onset of viscous thinning (decrease of $G^{\prime\prime}_{1}$) occurs when the contribution of $v_{3}$ is still positive. This could be due to the non-constant shear rate applied in PP geometries that leads to different microstructural responses along the geometry radius, thus having no significant microstructural relevance, or it could signify a momentary strain-dependence of the non-linearity. Considering the second hypothesis, due to the time-dependent nature of oscillatory shear, it could be possible that for a small strain, chain rotation is sufficient to comply with the applied deformation, independently of the strain rate, given that the oscillatory frequency is small. As such, only beyond a critical strain would the microstructure begin to breakdown due to viscous shearing forces. We performed a short numerical work to understand the mechanism behind the beginning of viscous thinning.
@@ 3.3 Numerical analysis
We set out to model the response of a magnetised-particle chain under oscillatory shear in COMSOL MultiPhysics, seeking to gather information on the critical strain required to provoke chain subdivision. The three-dimensional rotational flow imposed by the rheometer was approximated by a two-dimensional flow in a straight rectangular channel ($h_{\mathrm{c}}\times 6h_{\mathrm{c}}$, with $h_{\mathrm{c}}=0.1$ mm) driven by a velocity of the top channel wall, $U$. The domain was meshed with 2400 equal-sized quadrilateral elements (of side $l=h_{\mathrm{c}}/20$)ivivivFor the sake of conciseness, we present a mesh analysis as supplementary material, Section S4. The flow was calculated by solving the equations of conservation of mass and momentum for the isothermal, incompressible and laminar flow of a Newtonian fluid:
$\rho\left(\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\mathbf{\nabla}\right)\mathbf{u}=-\mathbf{\nabla}p+\eta\mathbf{\nabla}^{2}\mathbf{u}+\mathbf{F}\,$, (10)
$\mathbf{\nabla}\cdot\mathbf{u}=0\,.$ (11)
No-slip boundary conditions were imposed at the top and bottom walls, and open-boundary conditions were applied at the lateral walls, allowing free flow in and out of the channel. Figure 8 presents a schematic of the numerical model.
Figure 8: Schematic of the numerical model displaying the employed mesh, boundary conditions and fluid velocity imposition at the top wall ($U$).
@@
$\boldsymbol{F_{\mathrm{d}}}=3\,\pi\,\eta\,d_{\mathrm{p}}(\boldsymbol{u_{\mathrm{c}}}-\boldsymbol{u_{\mathrm{p}}})\,$, (12)
$\boldsymbol{F_{\mathrm{a,ij}}}=\frac{3\,\mu_{0}\,\mu_{\mathrm{cr}}\,m^{2}}{4\,\pi\,r_{\mathrm{ij}}^{4}}\left[\left(1-5\left(\hat{m}\cdot\hat{r}_{\mathrm{ij}}\right)^{2}\right)\hat{r}_{\mathrm{ij}}+2\left(\hat{m}\cdot\hat{r}_{\mathrm{ij}}\right)\hat{m}\right$]\,, (13)
$\boldsymbol{F_{\mathrm{r,ij}}}=2\,\frac{3\,\mu_{0}\,\mu_{\mathrm{cr}}\,m^{2}}{4\,\pi\,d_{\mathrm{p}}^{4}}\exp\left[{-30\left(\frac{r_{\mathrm{ij}}}{d_{\mathrm{p}}}-1\right)}\right]\hat{r}_{\mathrm{ij}}\,$, (14)
where $(\boldsymbol{u_{\mathrm{c}}}-\boldsymbol{u_{\mathrm{p}}})$ is the particle velocity ($\boldsymbol{u_{\mathrm{p}}}$) relative to the continuous phase ($\boldsymbol{u_{\mathrm{c}}}$), $\mu_{0}$ is the vacuum permeability, $\mu_{\mathrm{cr}}$ is continuous phase relative permeability ($\mu_{\mathrm{cr}}=1$), $\boldsymbol{r_{\mathrm{ij}}}$ is the interparticle vector (magnitude $r_{\mathrm{ij}}$ and direction $\hat{r}_{\mathrm{ij}}$) and $\boldsymbol{m_{\mathrm{ij}}}$ is the magnetic moment (with magnitude $m=4/3\pi(d_{\mathrm{p}}/2)^{3}\,\rho_{\mathrm{p}}\,M^{*}$ and direction $\hat{m}_{\mathrm{ij}}$, perpendicular to the flow direction). In Figure 9, a schematic of the forces applied to a representative 2-particle chain is shown. Regarding the temporal discretization, we employed an adaptive scheme, allowing COMSOL to adjust the time step freely, except for the initial value of the particle simulation, which was defined very small to guaranty convergence ($\Delta t_{\mathrm{0}}=10^{-6}$ s).
@@
Before proceeding with the oscillatory shear simulations, we tested for steady shear first to evaluate the critical shear rates for chain subdivision and complete breakdown. The shear rate was varied by imposing a top wall velocity $U=\dot{\gamma}\,h_{\mathrm{c}}$. It was found that both chains breakdown to individual particles when $\dot{\gamma}_{\mathrm{brk.}}\approx 0.79$ s-1, which was expected because chain breakdown is equivalent to the subdivision of any two-element chain. On the other hand, for $l_{\mathrm{chain}}=4$ the critical subdivision shear rate was $\dot{\gamma}_{\mathrm{sub.}}\approx 0.23$ s-1, while the largest chain with $l_{\mathrm{chain}}=10$ subdivided earlier, at $\dot{\gamma}_{\mathrm{sub.}}\approx 0.04$ s-1 (data not shown here for conciseness). Still considering steady shear, we can also gather information on the critical angle for chain break, $\phi_{\mathrm{c}}$. It is relevant, however, that while in reality the particle contact force is a step function triggered for $r_{\mathrm{ij}}=d_{\mathrm{p}}$, the excluded-volume force in our formulation is significant for small interparticle distances and a global magnetic force can be computed as $\boldsymbol{F_{\mathrm{m,ij}}}=\boldsymbol{F_{\mathrm{a,ij}}}+\boldsymbol{F_{\mathrm{r,ij}}}$. In Figure 10 are plotted the forces acting on a 2-element chain subject to a steady flow with shear rate $\dot{\gamma}=1.1$ and 0.7 s-1 against the chain angle, $\phi$. On the left column of Figure 10 are shown the global magnetic and viscous drag forces, and on the right column, the magnetic attraction and repulsive excluded-volume forces are shown decoupled (see Figure 9 for reference).
@@
We can see that the viscous shearing force weakens slightly as the chain rotates due to the reduction of the vertical disparity between the particles ($\Delta y=r_{\mathrm{ij}}\,\cos(\phi)$). Contrarily, the magnetic force increases on both axes ($x$ and $y$) as the magnetic dipole interactions try to re-align the chain with the magnetic field (aligned with $y$). If this simultaneous decrease of the relative drag and increase of the magnetic force is sufficient to reach an equilibrium ($F_{\mathrm{d,x}}+F_{\mathrm{m,x}}=0$) the chain remains stable, as is the case for $\dot{\gamma}=0.7$ s-1. However, if, despite the force convergence, the viscous drag remains too strong (the case where $\dot{\gamma}=1.1$ s-1), the chain breaks. Computing the minimum of the force difference $F_{\mathrm{d,x}}+F_{\mathrm{m,x}}$, we reach the critical chain angle $\phi_{\mathrm{c}}\approx 33.5^{\circ}$ (plotted as a vertical dotted line). With the current method, however, the critical angle is slightly dependent on the applied shear rate. This is due to the excluded volume force feeding instantaneous distancing that arises from the viscous effects, thus inhibiting pure rotation and contributing to premature chain break. As can be seen from the plots of the decoupled magnetic attraction, the excluded-volume force reduces the stability of the chain. Still, it also delays the critical angle, as the minimum of the function $F_{\mathrm{d,x}}+F_{\mathrm{a,x}}$ is found slightly earlier, at $\phi_{\mathrm{c}}\approx 28.3^{\circ}$.